A weather balloon rises through the atmosphere, its volume expanding from 3.4 m3 to 14 m3 as the temperature drops from 22 ∘C to -10 ∘C. If the initial gas pressure inside the balloon is 1.0 atm, what is the final pressure?
Answer #1

Ans=  V1=3.4m^3
V2=14m^3
t1=22 C=273+22=295K
t2=-10.0 C=273-10=263K
p1=1 atm= 10^5N/m^2
p2=?

p1*V1/T1=p2*V2/T2 =>
p2=p1*V1*T2/T1*V2
= 10^5*3.4m^3*263K/295*14
=21651.33172 N/m^2 =0.216 atm